Circular No. 6413

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

EXO 2030+375 M. T. Stollberg, University of Alabama in Huntsville; R. B. Wilson,

Marshall Space Flight Center; M. H. Finger, Universities Space Research Association; and T. A. Prince, California Institute of Technology, report for the Compton Gamma Ray Observatory BATSE team: “Pulsations from the Be/x-ray transient pulsar EXO 2030+375 have been observed during Apr. 1–7 and May 17–22. Outbursts began 4–5 days before the periastron passages predicted by the orbital model of Stollberg et al. (1993, AIP Conference Proc. 308, 255). All prior outbursts observed by BATSE occurred after periastron. The measured pulse period for the first outburst, after correcting for orbital motion, was 41.6773 ± 0.0015 s on 1996 Apr. 2.5 UT. The rms pulsed flux (20–50 keV) peaked on that day at 16 ± 3 mCrab. The second outburst had an orbit-corrected pulse period of 41.67781 ± 0.00083 s on 1996 May 19.5. The rms pulsed flux (20–50 keV) peaked at 15 ± 3 mCrab on May 21.5. Thirteen consecutive outbursts were observed previously by BATSE from 1992 Feb. 8 to 1993 Aug. 28. During that time, the pulse period of EXO 2030+375 decreased from 41.6839 ± 0.0011 s on 1992 Feb. 10.5 to 41.6708 ± 0.0015 s on 1993 Aug. 23.5 for a mean period rate of −8.5 ms/yr. Over the 21 orbital periods during which the source was not detected, the mean period rate was +2.3 ms/yr. The next periastron passage is predicted to occur on 1996 July 7.”

X-RAY DETECTIONS OF COMETS K. Dennerl, J. Englhauser, and J. Trümper, Max-Planck-Institut für

Extraterrestrische Physik, Garching, communicate: “X-ray emission from several comets has been detected in data obtained with the Position Sensitive Proportional Counter (PSPC) of ROSAT during the all-sky survey, in addition to the discovery of x-rays from comet C/1990 N1 (IAUC 6404). We list here all of the significant comet detections obtained so far, with the corresponding all-sky survey observing intervals: C/1990 K1 (Levy), 1990 Sept. 6.50–7.10 UT, 1991 Jan. 10.82–11.96, 1991 Jan. 17.49–17.96; C/1990 N1 (Tsuchiya-Kiuchi), 1990 Nov. 18.76–20.36, 1991 Jan. 12.07– 13.67; 45P/Honda-Mrkos-Pajdušáková, 1990 July 31.20–31.86. All comets were at heliocentric distances < 2 AU, with visual m_{1} between 4.2 and 11.6. For a comparison of the x-ray and optical properties, we are looking for images of these comets taken at or near the specified times; please contact K.

Dennerl (e-mail kod@mpe-garching.mpg.de).”

1996 June 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green